### v1

# Context-Aware Code Wiring Recommendation with LLM-based Agent

Abstract
Copy-paste-modify is a widespread and pragmatic practice in software development, where developers adapt reused code snippets—sourced from platforms such as Stack Overflow, GitHub, or LLM outputs—into their local codebase. A critical yet underexplored aspect of this adaptation is code wiring, which involves substituting unresolved variables in the pasted code with suitable ones from the surrounding context. Existing solutions either rely on heuristic rules or historical templates, often failing to effectively utilize contextual information, despite studies showing that over half of adaptation cases are context-dependent. In this paper, we introduce WIRL, an LLM-based agent for code wiring framed as a Retrieval-Augmented Generation (RAG) infilling task. WIRL combines an LLM, a customized toolkit, and an orchestration module to identify unresolved variables, retrieve context, and perform context-aware substitutions. To balance efficiency and autonomy, the agent adopts a mixed strategy: deterministic rule-based steps for common patterns, and a state-machine-guided decision process for intelligent exploration. We evaluate WIRL on a carefully curated, high-quality dataset consisting of real-world code adaptation scenarios. Our approach achieves an exact match precision of 91.7% and a recall of 90.0%, outperforming advanced LLMs by 22.6 and 13.7 percentage points in precision and recall, respectively, and surpassing IntelliJ IDEA by 54.3 and 49.9 percentage points. These results underscore its practical utility, particularly in contexts with complex variable dependencies or multiple unresolved variables. We believe WIRL paves the way for more intelligent and context-aware developer assistance in modern IDEs.
Index Terms:
Copy-paste-modify Practices, Code Reuse, Code Adaptation, Large Language Models, Agent.
I Introduction
Copy-paste-modify is a common and inevitable practice during development. The developers frequently reuse code snippets from online programming Q&A communities, e.g., Stack Overflow [16], open-source repositories, e.g., GitHub [15], or code generation of LLMs. However, in most cases the reused code snippets need additional adaptations to integrate to the local codebase. As reported, more than 85% code snippets from Stack Overflow need to be adapted (modified) before integration into the local code [46, 47]. Variable identifiers are usually under adaptation since undeclared or conflict identifiers would lead to compilation errors or potential bugs. Consequently, code wiring, replacing unresolved variables with existing ones from the local context, is one of the most prevalent forms of adaptation. This term aligns with the analogy of electrical wiring, where components (variables in code) are intentionally connected to establish functionality. An example of code wiring from Stack overflow [9] to GitHub [10] is presented in Fig. 1. Automatic code wiring can help developers get rid of the repetitive and trivial processes during the adaptation and concentrate on the complex business logic. However, existing approaches are designed based on either simple heuristic rules [6] or templates extracted from historical modifications [12], leaving the context not effectively leveraged although 56.1% of the adaptation cases are dependent on the surrounding context as reported by Zhang et al. [46]. Full-parameter LLMs have demonstrated promising performance; however, their high latency renders them impractical for real-world applications. On the other hand, distilled LLMs with smaller parameter sizes tend to under perform in terms of accuracy and robustness.
To this end, we present WIRL, an LLM-based agent for code Wiring through Infilling with RAG and LLMs. WIRL consists of three core components: an LLM, a customized toolkit, and an agent pilot. The customized toolkit offers three primary functionalities: (1) identifying and locating unresolved elements, (2) analyzing and collecting contextual information, and (3) infilling and recommending suitable substitutions. The agent pilot is responsible not only for initializing the prompt but also for coordinating the interaction between the LLM agent and the toolkit. It parses the LLM’s output, updates the prompt dynamically, and invokes the appropriate tools as needed. With a dynamically updated prompt, WIRL incrementally gathers relevant context by invoking appropriate tools until the agent determines that sufficient information has been collected to make a final recommendation. To fully leverage the capabilities of LLMs, we reformulate the adaptation task as a retrieval-augmented generation (RAG)-based infilling task for the unresolved elements—framing it as a code completion problem, which aligns more naturally with the strengths of LLMs. Empirical evidence from Zhang et al. [30] supports this reformulation, showing that LLM performance on code snippet adaptation tasks lags behind their performance on code completion and generation tasks. Moreover, by adopting the RAG-based infilling strategy, even distilled LLMs with smaller parameter sizes can outperform full-parameter LLMs, thereby meeting both the latency and accuracy requirements of real-world software development scenarios. To enhance the efficiency of WIRL, we adopt a hybrid execution mode for the agent, wherein essential steps are extracted and executed initially without invoking the LLM. In addition to this optimization, we introduce a state machine to guide the agent’s exploration process, thereby reducing unnecessary or ineffective attempts and improving overall execution efficiency.
To assess the efficiency of WIRL, we manually curated a high-quality code wiring dataset derived from a publicly available code adaptation dataset [14]. Evaluated on this dataset, WIRL achieves an exact match precision of 91.7% and a recall of 90.0%. Notably, it outperforms advanced LLM baselines by 22.6 and 13.7 percentage points in precision and recall, respectively, and exceeds the performance of IntelliJ IDEA by 54.3 and 49.9 percentage points. These results highlight the effectiveness and practical advantage of WIRL. We also analyzed time efficiency and token consumption, demonstrating that WIRL meets real-time requirements with affordable computational cost.
The contributions of this paper are as follows:
  • We propose WIRL, an LLM-based agent specifically designed for context-aware code wiring recommendations.
  • We develop a customized toolkit that provides essential functionalities, enabling the agent to perform precise and efficient code adaptation.
  • We construct and release a high-quality dataset for code wiring evaluation, curated and validated through careful manual inspection.
The rest of this paper is structured as follows. Section II motivates this study. Section III illustrates the design of WIRL. Section IV presents experimental details and results. Section V discusses the threats and limitations of our study. Section VI discusses the related works, and Section VII concludes this paper.
II Motivating Example
To provide the intuition about how WIRL works and to define the key terminologies, we begin with a motivating example of a code wiring instance. Fig. 1 presents an example of copy-paste-modify (specifically code wiring) practice where a developer reused a piece of code snippet from Stack Overflow [9] and adapted it to the GitHub local repository [10]. The developer copied and pasted the line 6 and line 7 and then modified them into line 8 and line 9. For the sake of clarity, in this paper we call the code snippets in line 6 and line 7 isolated code since it is isolated before integration. The code snippets in line 8 and line 9 are called integrated code, i.e., code after the integration. Specifically, in this example, local variable “list” (highlighted in red) is undefined and leads to syntax errors. The developers substituted it with the predefined field variable “mTags”(highlighted in yellow and green) to resolve the syntax errors. We call “list” an unresolved element. It is worth noting that variables are not the only code entity that need to be modified, literal values or a method invocation expression could also be the target code entity. Consequently, we use “element” instead of only “variable”. “mTags” is called the context element, i.e., the code elements in the context. Similarly, a context element could also be either a variable (including local variables, parameters, and fields) or an expression (e.g., method invocation).
Before illustrating how WIRL operates, we first present the problem formulation, which serves as an important foundation of our approach. Assume the set of unresolved elements is denoted by $\mathcal{U}={u_{i}}$ and the set of context elements by $\mathcal{C}={c_{i}}$. The code wiring task can be formally defined as identifying an injective mapping $\mathcal{M}:\mathcal{U}\rightarrow\mathcal{C}$, where each $u_{i}\in\mathcal{U}$ is mapped to a distinct $c_{i}\in\mathcal{C}$. In this work, however, we reformulate code wiring as a retrieval-augmented generation (RAG)-based infilling task. Specifically, each $u_{i}\in\mathcal{U}$ is replaced with a placeholder token “<Infill>” in the isolated code. The objective then becomes to infill these placeholders using the retrieved context information, including the variable names themselves, which is motivated by the observation that literal similarity often provides valuable contextual cues. This reformulation is grounded in the insight that code completion is more naturally aligned with the capabilities of large language models (LLMs) than explicit code adaptation. This design choice is further supported by recent empirical findings by Zhang et al. [30], which highlight the superior performance of LLMs in code completion tasks relative to direct adaptation scenarios.
In this example, WIRL first identifies the unresolved variable “list”, and then proceeds to iteratively collect relevant contextual information such as available variables in the current scope, semantic hints indicating that “list” should refer to a collection object, and method names suggesting that “Tags” is the intended target. Then it assesses the suitability of candidate variables by invoking a set of external tools tailored for context analysis. Once the context is deemed sufficient, WIRL conduct infilling and “mTags” is filled as the appropriate variable here. Eventually, WIRL automatically applies the modification within the IDE for developers.
Figure 1: Motivating Example
III Approach
III-A Overview
Figure 2: Overview of WIRL
The overview of WIRL is illustrated in Fig. 2. WIRL consists of three core components: an LLM, a customized toolkit containing three types of tools, i.e., locator, collector, and completer, and an agent pilot, which orchestrates communication between the LLM and the toolkit by parsing outputs, updating prompts, and executing tool invocations. Given a piece of isolated code, the agent pilot begins by initializing the prompt (Step 1). This prompt includes system roles, task descriptions, and instructions on how the task should be addressed using the available tools. It also embeds the location of unresolved elements and results from preliminary analyses. Using this prompt, the agent pilot issues a request to the LLM (Step 2). The LLM responds by specifying a tool to invoke from the customized toolkit (Step 3). The agent pilot then parses the response and executes the corresponding tool with the appropriate arguments (Step 4). The tool’s output is subsequently parsed and integrated into the prompt (Step 5) for the next iteration. This loop continues iteratively until either the completer tool is executed, indicating task completion, or a predefined computational budget is reached.
III-B Dynamic Prompting
WIRL adopts a reasoning-and-acting framework (i.e., ReAct [45]), iterating until either the goal is achieved or the computational budget is exhausted. In each iteration, the agent pilot parses the output of the invoked tool and incorporates the results into the prompt, which then serves as the input for the subsequent iteration. The static and dynamic sections of the prompt are described in detail below:
III-B1 Role (Static)
The static section defines the agent’s role as a Senior Java Development Engineer and clearly articulates the task objective: to perform variable-level code integration based on the provided code context. Notably, this section explicitly instructs the agent to make decisions autonomously, without relying on user intervention at any point during the process.
III-B2 Goals (Static)
We define three primary goals for the agent to accomplish:
  • Analyze context: Evaluate the current code context to determine whether it provides sufficient information to suggest a valid substitution.
  • Collect more context: If the existing context is inadequate, invoke the appropriate tools from the provided toolkit to gather additional relevant information.
  • Suggest a substitution: Once adequate context is obtained, leverage the agent’s completion tool to recommend a suitable substitution for the unresolved element.
III-B3 Guidelines (Static)
We define a set of guidelines to which the agent should adhere.
  • Limited budget: We emphasize that the budget is limited and it is important to be efficient to select the next steps.
  • Type should match: We inform the agent that the basic principle is to make sure the retrieved context element is type-compatible with unresolved element.
  • Be careful to stretch: More context typically leads to higher latency. Consequently, the agent is instructed to complete its recommendation using the minimal necessary context and to extend the context scope cautiously.
  • Extend scope: We instruct the agent that in some cases, no suitable substitution candidates may be found within the immediate context. In such scenarios, the agent should expand its search scope beyond the current class, as the context element is more likely to be a method invocation expression rather than a simple variable.
III-B4 State Description (Dynamic)
To guide the agent in utilizing the customized toolkit in a meaningful and efficient manner, we define a state machine comprising three states: initial state, insufficient context state, and sufficient context state. This mechanism is introduced because, in our preliminary experiments, we observed that the agent often engaged in aimless exploration when operating without explicit guidance.
Each state is associated with a set of available tools, as described in Section III-C, and the state description section of the prompt informs the agent of its current status. Specifically, the agent begins in the initial state, where it identifies unresolved elements and performs a preliminary analysis of the context. Based on this analysis and subsequent tool executions, the agent transitions into the insufficient context state. In this state, it iteratively invokes relevant tools to gather additional contextual information until either the context is deemed sufficient or the predefined resource budget is exhausted. Once the agent reaches the sufficient context state, it invokes the completion module (also implemented as an LLM invocation, detailed in Section III-C) to make the final recommendation.
It is important to note that while the state machine offers structured guidance, it does not enforce a fixed sequence of tool invocations. The selection and ordering of tools are left entirely to the agent’s discretion.
III-B5 Available Tools (Dynamic)
This section illustrates the available tools (refer Section III-C) agent can call in each state.
III-B6 Gathered Information (Dynamic)
A key capability of WIRL lies in its ability to gather contextual information through tool execution, which forms the foundation for generating accurate recommendations. To ensure the agent remains aware of previously executed tools and the context collected, this section of the prompt records the reasoning thoughts, the name of tool, its corresponding arguments (parsed from the LLM’s output), along with the results returned by tool invocations (parsed from the tools’ output).
III-B7 Output Format (Static)
To enable the agent pilot to effectively parse the output, we require that all responses from the LLM be structured in JSON format. Each JSON object must contain three mandatory fields: “thought”, “action”, and “action_input”. The “thought” field captures the agent’s reasoning and decision-making process based on the current context. The “action” field specifies the name of the tool to be invoked next, while the “action_input” field provides the necessary arguments for that tool, also formatted as a JSON object.
III-C Customized Toolkit
TABLE I: Available Tools for the Agent
Tools Type Applicable State Description
identify_unresolved_elements Locator Initialization Analyze compiler information to identify the unresolved elements.
get_available_variables Collector Initialization Conduct code analysis and get the available variables in the current context scope.
get_unused_variables Collector Initialization Conduct data flow analysis and get the unused variables in the current context.
is_argument Collector Initialization Judge whether the unresolved element plays an argument in a method invocation.
is_receiver Collector Initialization Judge whether the unresolved element plays a receiver in a method invocation.
retrieve_identical_function_call Collector Insufficient Context Retrieve variables that invoke the identical function calls with unresolved element.
reserve_type_compatible_ones Collector Insufficient Context Only reserve the variables that are type-compatible as the unresolved element.
sort_by_literal_similarity Collector Insufficient Context Sort the available variables by their similarity with the unresolved element.
get_method_names Collector Insufficient Context Collect method members with same type as unresolved element in the given class.
execute_completion Completer Sufficient Context Invoke LLM to complete the code snippets with the collected context information.
One of the key novelties of WIRL lies in its ability to autonomously decide which tools to invoke based on the current state. The toolkit provided for WIRL (as shown in Table I) is carefully customized to facilitate the code wiring task.
III-C1 Identifying unresolved elements
A prerequisite for effective code wiring is the accurate identification of unresolved elements. The locator tool (identify_unresolved_elements) leverages compiler information to detect unresolved variables and their references, and it is executed as the initial step. Since identifying and locating these elements is fundamental to addressing the task, the agent pilot invokes the locator automatically, without requiring input from the LLM.
III-C2 Collecting Context Information
To minimize unnecessary LLM invocations and reduce iteration overhead, we provide two tools, i.e., get_available_variables and get_unused_variables, to help the agent efficiently construct a list of candidate variables. The get_available_variables tool parses the AST of the current Java file and collects all accessible local variables, method parameters, and class fields. To reflect realistic development scenarios, only local variables declared before the unresolved element are considered valid candidates. Once available variables are gathered, the get_unused_variables tool performs data-flow analysis to examine the usage of each variable. Variables with no detected references are marked as unused. To determine the syntactic role of an unresolved element, the tools is_argument and is_receiver inspect its AST context to identify whether it is enclosed within a method invocation. If the element appears as a method argument, its expected type can be inferred from the corresponding formal parameter, and the parameter name itself provides useful semantic cues. If the element is the receiver of a method call, the tool returns the invoked method member to support further analysis. The reserve_type_compatible_ones tool filters candidate variables by retaining only those whose types are compatible with that of the unresolved element. When the unresolved element is identified as a method receiver, the retrieve_identical_function_call tool takes the invoked method member as input and searches the current file for instances of the same member invocation, which may reveal contextually relevant patterns or variable usages. Finally, if no valid candidates are found in the current scope, the get_method_names tool is invoked. Given a class name, it returns the method members of that class that match the type of the unresolved element, providing additional opportunities for accurate substitution. The sort_by_literal_similarity tool is invoked when the agent determines that lexical similarity is a relevant factor. This tool computes the Levenshtein distance between the unresolved element and each candidate variable, ranking the candidates based on their literal similarity.
III-C3 Infilling Isolated Code
Once the agent has collected sufficient contextual information, it invokes the execution_completion tool to infill the placeholders using the gathered data. This tool is implemented via an LLM call, where the prompt corresponds to the final version of the dynamically updated prompt that incorporates all previously collected context information.
III-D Agent Pilot
The agent pilot orchestrating the communication between LLM and customized toolkit, playing a essential role in WIRL. Its responsibilities are outlined as follows:
III-D1 Initializing Prompt
The agent pilot begins by initializing the prompt with static components such as the role definitions and task objectives. It then incorporates the input code (as described in Section IV-E1) into the prompt. Subsequently, the prompt is dynamically updated with the locations of unresolved elements and relevant contextual information, obtained by invoking the Locator (identify_unresolved_elements) and Collector tools (get_available_variables, get_unused_variables, is_argument, and is_receiver). This initialization process substantially reduces unnecessary LLM invocations and accelerates context analysis, thereby enhancing the overall efficiency of WIRL.
III-D2 Parsing Output
Since the outputs of LLMs are returned in JSON format, they can not be directly used for tool execution. Accordingly, the second responsibility of the agent pilot is to validate and process these outputs. If an LLM response deviates from the expected structure, e.g., due to hallucinations or formatting errors, the agent pilot must attempt to correct the output or handle exceptions gracefully. In addition to output parsing and error handling, the agent pilot also maintains a memory of previously executed tools and their corresponding arguments to prevent redundant operations.
III-D3 Executing Tools
With the parsed tool names and corresponding arguments, the agent pilot invokes the designated tools and returns their outputs. Notably, all tool executions are performed within an isolated environment to ensure they do not interfere with the host system or the internal functioning of WIRL.
III-D4 Updating Prompt
Once the tool output is available, the agent pilot updates all dynamic sections of the prompt in preparation for the next iteration. Specifically, it modifies the current state and the list of available tools, and appends the “thought”, “action”, “action_input”, and “observation” (i.e., the tool’s output) to the accumulated context information.
IV Evaluation
IV-A Research Questions
  • RQ1: How well does WIRL perform compared with baselines?
  • RQ2: How well does WIRL perform regarding time cost?
  • RQ3: How well does WIRL perform regarding token consumption and monetary cost?
RQ1 investigates the effectiveness of WIRL in comparison to selected baseline approaches for automatic code wiring. By addressing this question, we aim to understand how well WIRL performs in real-world development scenarios, as the evaluation leverages real data under realistic experimental conditions. Additionally, this analysis helps identify specific cases where WIRL outperforms the baselines. Gaining insights into the strengths of WIRL can guide future improvements and optimizations of the approach.
RQ2 examines the time efficiency of WIRL relative to baseline methods, with a particular focus on whether WIRL can meet the responsiveness requirements of integrated development environments (IDEs) and developers. To answer this question, we use the average time taken to complete a code wiring task as the primary metric. This allows us to assess whether WIRL delivers timely recommendations that support fluid and productive software development workflows.
RQ3 assesses the token consumption and associated monetary cost of WIRL. Specifically, we evaluate input/output token usage and the total cost incurred per code wiring instance. By comparing these metrics against those of the baseline approaches, we aim to determine the cost-efficiency of WIRL in delivering practical code wiring support at a reasonable computational expense.
IV-B Dataset
To the best of our knowledge, there are no existing datasets specifically collected for the evaluation of automatic code wiring approaches. Recently, Zhang et al.[46] investigated context-based code snippet adaptation and constructed a high-quality dataset through a combination of automated processing and meticulous manual curation. The resulting dataset comprises 3,628 real-world code reuse cases, where code snippets were copied from Stack Overflow and subsequently adapted for use in GitHub projects. Their dataset was initially derived from the latest version of SOTorrent[41] (version 2020-12-31), available via Zenodo [17], and includes only those reuse cases in which GitHub files contain explicit references to Stack Overflow posts. Given the quality and relevance of this dataset, we adopted it as the foundation for our evaluation and performed additional filtering and manual curation. The construction of the evaluation dataset used in this paper involves the following steps:
  • To prevent data leakage and potential overfitting, we first excluded the 300 sampled instances used by Zhang et al. [46], as the design of WIRL was partially inspired by their empirical observations.
  • From the remaining 3,328 reuse cases, we performed a deduplication step to eliminate redundancy caused by forked repositories. Subsequently, we constructed a mapping between GitHub repositories and the corresponding Stack Overflow posts.
  • As WIRL depends on compiler feedback and AST binding analysis, it requires the full project to be resolvable. Therefore, we conducted a project-wise manual inspection of the remaining dataset. The mappings were sorted in descending order based on the frequency of references to each repository, prioritizing more frequently reused code.
  • Following this order, two authors independently reviewed each code reuse case to determine whether it qualifies as a code wiring instance. In cases of disagreement, discussions were held until a consensus was reached. Ultimately, we identified 100 code wiring cases involved with 221 pairs of unresolved elements and context elements from Stack Overflow to GitHub. For ease of reference, we refer to this curated evaluation dataset as $CWEvaluation$.
IV-C Selected Baselines
IV-C1 SOTA Approaches
We selected ExampleStack [28] as a baseline because it represents the state-of-the-art in automatic code wiring. ExampleStack is a Google Chrome extension designed to assist developers in adapting and integrating online code examples into their own code repositories. The authors provide comprehensive reproduction instructions, which allowed us to successfully set up and run ExampleStack in our evaluation.
Besides that, we selected the widely adopted industrial IDE, IntelliJ IDEA [6], as one of our baseline approaches. The rename recommendation support in IDEA is both convenient and practical, and it is also suitable to address the code wiring problem.
IV-C2 LLMs
We select the following LLMs as our baselines: Distilled tiny models, including GPT-4o-mini [2], Qwen2.5-Coder-14B [3], and Qwen2.5-Coder-32B [4]. Full-parameter models, including DeepSeek-V3 [1] and Qwen-max [5].
IV-D Metrics
We adopted the evaluation metrics used by Wang et al. [44] to assess the performance of WIRL and the baselines. The definitions are introduced in the following:
  • #Total Cases: the number of cases involved in the evaluation, i.e., number of unresolved elements.
  • #Recommendation (Rec): the number of cases where the evaluated approaches make a recommendation for the developers. Note that the number of recommendations (#recommendation) may not always match the total number of cases (#total cases).
  • #Exact Match (EM): the number of cases where the recommended code elements are identical to the ground truth (i.e., the existing names in context).
  • $EM_{Precision}$: the number of exact matches divided by the number of recommendations, i.e.,
    $EM_{Precision}=\frac{\#Exact\ Match}{\#Recommendation}$ (1)
  • $EM_{Recall}$: the number of exact matches divided by the number of cases involved in the evaluation.
    $EM_{Recall}=\frac{\#Exact\ Match}{\#Total\ Cases}$ (2)
IV-E Experimental Setup
IV-E1 Input and Output Format
To better simulate real-world development scenarios, the input code format is configured as follows:
  • For the LLM baselines, the entire class containing the isolated code is provided as input. The isolated code, representing the segment requiring adaptation, is explicitly marked using a pair of control tokens, $<start>$ and $<end>$, to delineate the adaptation region. The complete prompt format used for the LLM baselines is available in our public repository [8].
  • To enhance efficiency, the input to WIRL consists solely of the method declaration containing the isolated code. WIRL is designed to operate with the minimal necessary context and selectively expand the context scope when needed.
  • Furthermore, we assume that any code following the isolated code is unavailable, reflecting the common top-down coding pattern observed during software development. In this scenario, only the code preceding the adaptation region, which within the same method declaration, is accessible as context.
For the sake of clarity, we constrained the LLM baselines to output only the unresolved elements along with their corresponding context elements. To promote self-consistency and mitigate output variability, we set the temperature to 0. Furthermore, each evaluated approach was executed five times, and the most frequently generated output was selected as the final result, following the protocol outlined by Chen et al. [18].
As both IDEA and ExampleStack produce ranked lists of candidate substitutions, we only considered their top-ranked recommendation as the final answer to ensure a fair and consistent comparison across all evaluated approaches.
IV-E2 Implementations
To implement WIRL, we built a plugin for IntelliJ IDEA using the IntelliJ PlatForm Plugin SDK [7] and the Langchain4J [13] framework to interface with LLM APIs. While WIRL is designed to be compatible with any advanced LLMs, in this study we use Qwen2.5-Coder-14B as the backend model for two key reasons. First, if WIRL demonstrates superior performance over state-of-the-art full-parameter LLMs while relying on a smaller model, it further highlights the strength of our design. Second, models with fewer parameters typically respond faster [43, 42], making them more suitable for real-time usage in development environments.
To support code analysis and contextual understanding, we leverage the Program Structure Interface (PSI) [11], a powerful abstraction for syntactic and semantic analysis, including data flow inspection and code structure extraction.
In addition, WIRL employs an iterative strategy to perform context-aware code wiring. To balance latency, computational cost, and effectiveness, we empirically set the maximum number of iterations at two.
IV-F RQ1: Outperform the SOTA
TABLE II: Comparison with Baselines
#EM #Rec $EM_{Precision}$ $EM_{Recall}$
IDEA 79 189 41.8% 35.7%
ExampleStack 4 10 40.0% 1.8%
GPT-4o-mini 101 145 69.7% 45.7%
Qwen2.5-Coder-14B 110 157 70.1% 49.8%
Qwen2.5-Coder-32B 149 197 75.6% 67.4%
Qwen-max 124 159 78.0% 56.1%
DeepSeek-V3 129 171 75.4% 58.4%
WIRL - Locator 110 157 70.1% 49.8%
WIRL - Collector 169 190 88.9% 76.5%
WIRL - Completer 146 189 77.2% 66.1%
WIRL 199 217 91.7% 90.0%
In Table II, the first column lists the evaluated approaches. The second and third columns indicate, respectively, the number of cases in which the recommended code elements exactly match the ground truth and the number of cases in which a recommendation is made by the approach. The final two columns, $EM_{Precision}$ and $EM_{Recall}$, report how often the recommendations are correct and how many of the correct substitutions are successfully identified, respectively (see Section IV-D for detailed definitions). From Table II, we make the following observations:
  • WIRL outperforms all selected baselines by a significant margin in both $EM_{Recall}$ and $EM_{Precision}$. It produces the highest number of recommendations (217) while also achieving the highest number of exact matches (199). The resulting $EM_{Recall}$ and $EM_{Precision}$ scores of 90.0% and 91.7%, respectively, reflect its strong and reliable performance in real-world development scenarios.
  • Compared to IDEA, WIRL achieves improvements of 54.3 percentage points in $EM_{Recall}$ and 49.9 percentage points in $EM_{Precision}$. Against ExampleStack, it shows even greater gains: 88.2 and 51.7 percentage points in $EM_{Recall}$ and $EM_{Precision}$, respectively. When benchmarked against the advanced LLM model Qwen2.5-Coder-32B, WIRL still delivers notable improvements of 22.6 and 13.7 percentage points, further demonstrating its superior effectiveness.
  • WIRL also substantially improves upon the performance of directly using a raw LLM (Qwen2.5-Coder-14B). Without the academic design described in Section III, the baseline LLM achieves only 49.8% $EM_{Recall}$ and 70.1% $EM_{Precision}$. With the integration of the external toolkit and structured prompt design, WIRL improves $EM_{Recall}$ and $EM_{Precision}$ by 40.2 and 21.6 percentage points, respectively, highlighting the value of its architectural innovations.
  • The three types of tools in WIRL contribute differently to its performance: locator tools have the greatest impact, followed by completer tools, with collector tools contributing the least. Locator tools contribute the most because the collector and completer tools rely on the location information provided by them. Removing the locator tools will reduce the system to relying solely on raw LLMs.
We further analyzed the reasons behind WIRL’s superior performance compared to baseline approaches, with a particular focus on IDEA. During the evaluation of IDEA, we recorded not only the top-ranked recommendation but also the complete list of suggestions. Our quantitative and qualitative analysis revealed that in 84.1% of the cases (159 out of 189), IDEA did include the correct recommendation, but it was not ranked first. Statistically, the correct answer appeared in the top position in only 49.7% of those 159 cases. Notably, in some instances, the recommendation list contained more than 100 items, requiring developers to scroll extensively before locating the correct option. These findings indicate that while IDEA’s underlying recommendation algorithm, originally designed for renaming suggestions, can potentially address the code wiring problem, it lacks an effective ranking mechanism to prioritize relevant suggestions. This inefficiency increases the cognitive load and selection time for developers. In contrast, WIRL generates a single high-confidence recommendation for each unresolved element, achieving a high $EM_{Precision}$ of 91.7%. This direct substitution approach not only reduces the time and effort required from developers but also allows them to focus more on adapting business logic rather than sifting through lengthy suggestion lists. Furthermore, IDEA failed to generate any recommendation in 14.5% of the cases (32 out of 221). These failures occurred when the unresolved elements were expressions rather than variables or literals, i.e., cases that exceed the intended scope of IDEA’s renaming functionality. In contrast, WIRL successfully handled such cases through its use of static code analysis, showcasing its broader applicability and robustness.
For ExampleStack, we conducted a manual analysis of the input data and identified the key limitation behind its unsatisfied performance. The core rationale of ExampleStack depends on a pre-collected database of historical modifications of code snippets from Stack Overflow to GitHub. Consequently, its effectiveness is heavily constrained by the presence or absence of identical or similar historical modifications in the database. This reliance significantly hampers its applicability to previously unseen cases, which are common in real-world development scenarios. In fact, within the $CWEvaluation$, only 6 code wiring instances involving 18 unresolved elements had similar modifications present in the original database provided by ExampleStack [12]. Across all test cases, ExampleStack made 10 recommendations, only 4 of which were correct. In contrast, WIRL leverages the reasoning and generation capabilities of large language models (LLMs) to recommend context elements dynamically and from scratch. This design enables WIRL to operate independently of historical reuse data, thereby overcoming the fundamental limitation of ExampleStack and improving its generalization and effectiveness across diverse code wiring scenarios.
Regarding the performance of raw LLMs, their relatively low $EM_{Recall}$ can be attributed to two main factors. First, the length of the input class significantly impacts the ability of LLMs to make accurate recommendations. When the class size exceeds the model’s maximum token limit, the LLM cannot process the full context, resulting in failure to generate any meaningful output. For instance, the maximum token limit for DeepSeek-V3 is 64K tokens. If a class exceeds this size, DeepSeek-V3 becomes incapable of handling the instance entirely. Second, raw LLMs sometimes fail to identify all unresolved elements within the code. As illustrated in Fig. 3, there are two unresolved elements, i.e., “list” and “target”, with their corresponding context elements being “listView” and “mCommentListPosition”. While Qwen2.5-Coder-14B successfully identified “list” and recommended “listView”, it failed to detect “target”, thus missing a key aspect of the task. In contrast, WIRL, by incorporating compiler information, accurately identified both unresolved elements (including all six references to “list”) and successfully recommended the appropriate context elements, “listView” and “mCommentListPosition”, by leveraging the reasoning capability of LLMs.
Figure 3: Example of Identification Failure for Raw LLM
The relatively low $EM_{Precision}$ of raw LLMs can be attributed to their difficulty in analyzing long and complex contexts to correctly identify mapping relationships, particularly when the appropriate context element is not explicitly present in the immediate code. As illustrated in Fig. 4, the unresolved element is encoding, and its corresponding context element is “Charset.defaultCharset()”. It is challenging for raw LLMs to recommend “Charset.defaultCharset()” without the context information that the “Charset” class provides a static method “defaultCharset()” that returns an instance of “Charset”. In contrast, WIRL addresses this limitation by leveraging its toolkit. The initial analysis of the context (is_argument tool) determines the type of encoding, which is “Charset”. Then, it uses the get_method_names tool to retrieve methods in the “Charset” class that return the same type. This process successfully identifies “defaultCharset()”, enabling WIRL to include it in the prompt for the final completion module, which then makes the correct recommendation. Additionally, raw LLMs occasionally generate substitutions beyond the unresolved elements, introducing unexpected recommendations and increasing the number of false positives, which further impacts $EM_{Precision}$.
Figure 4: Example of Recommendation Failure for Raw LLM
IV-G RQ2: Time Efficiency
TABLE III: Time, Token Consumption, and Monetary Cost
Avg. Time Cost (ms) Avg. #Input Tokens Avg. #Output Tokens Avg. #Total Tokens Avg. Monetary Cost (CNY)
IDEA 3.7 - - - -
ExampleStack 729.1 - - - -
GPT-4o-mini 3,930.4 2,769.4 27.4 2,796.8 0.0031
Qwen2.5-Coder-14B 4,624.5 2,769.4 28.3 2,797.7 0.0057
Qwen2.5-Coder-32B 6,799.4 2,769.4 43.8 2,813.2 0.0058
Qwen-max 5,393.6 2,769.4 24.9 2,794.3 0.0069
DeepSeek-V3 11,489.9 2,769.4 22.0 2,791.4 0.0057
WIRL 5,255.6 3,006.7 108.5 3,115.2 0.0067
In this research question, we examine the time efficiency of WIRL compared to other baseline approaches on the $CWEvaluation$. The comparison results are summarized in Table III, where the second column reports the average execution time per code wiring instance. It is worth noting that the reported time costs are averaged over five independent runs to ensure robustness and consistency. From Table III, we make the following observations:
  • WIRL required only marginally more time (approximately 0.6 seconds) to generate a recommendation compared to Qwen2.5-Coder-14B, and it demonstrated greater time efficiency than all other evaluated LLMs except GPT-4o-mini. This efficiency can be attributed to its use of a relatively compact LLM (14B parameters), whereas other models incur longer response times due to their significantly larger parameter sizes. For example, DeepSeek-V3, with 671B parameters, exhibited the highest latency, averaging 11.5 seconds per recommendation.
  • Although IDEA and ExampleStack achieved superior performance in terms of time cost, they fall short in ensuring recommendation accuracy. We argue that WIRL offers a more practical trade-off between recommendation quality and acceptable latency, making it well-suited for real-world development workflows.
To gain deeper insights into the time cost distribution of WIRL and the baseline approaches, we visualized the results using a violin plot, as shown in Fig. 5. The plot reveals that WIRL is capable of handling the majority of code wiring instances in under 5 seconds. Its median time cost is 4,201.5ms, which is comparable to that of Qwen2.5-Coder-14B (3,977.5ms), and substantially lower than that of DeepSeek-V3 (7,983.5ms). This demonstrates WIRL’s ability to deliver high-quality recommendations with competitive latency.
Figure 5: Time Cost Distribution
IV-H RQ3: Token Consumption and Monetary Cost
In this research question, we evaluate the token usage and monetary cost of WIRL compared to other LLM-based baselines on the $CWEvaluation$. The comparison results are summarized in Table III. Columns three to five report the average number of input tokens, output tokens, and total tokens per code wiring instance, respectively, while the final column shows the corresponding average monetary cost. All figures are averaged over five independent experimental runs.
As shown in Table III, WIRL consumes slightly more tokens than Qwen2.5-Coder-14B, with an average increase of 237.3 input tokens and 80.2 output tokens. This results in a marginally higher monetary cost, i.e., approximately 0.001 CNY more per instance. Surprisingly, although GPT-4o-mini incurs the lowest monetary cost, its performance is relatively poor, as shown in Table II. The increase in token usage is attributed to WIRL’s iterative design, where it progressively collects and integrates contextual information into the prompt before generating a final recommendation. As the output from one iteration is passed as input to the next, the cumulative token count naturally exceeds that of the single-pass LLM baselines.
To further examine the distribution of token consumption for WIRL and the LLM baselines, we present a violin plot in Fig. 6. The results indicate that WIRL handles the majority of cases using fewer than 2,500 tokens, with a median token count of 2,306.5 per code wiring instance, which is comparable to that of Qwen2.5-Coder-14B having a median of 1,962.5 tokens. Additionally, both WIRL and Qwen-max exhibit stable token usage, with minimal outlier points. Notably, Qwen-max shows consistent performance in both time and token consumption, largely because it omits cases that exceed its maximum token limit.
Figure 6: Token Number Distribution
V Discussion
V-A Threats to Validity
A potential threat to external validity lies in the limited size and diversity of the dataset used for evaluation. To mitigate this risk, we leveraged a publicly available dataset and performed a thorough manual inspection to ensure its quality. Additionally, we have open-sourced both the dataset and the evaluation results [8] to facilitate replication and further validation by the research community.
A threat to construct validity arises from possible bias in the manual identification of real-world code wiring instances. To address this, both the first and second authors independently inspected the data. The inter-rater agreement was high, with a Cohen’s kappa coefficient of 0.81, indicating strong consistency. Discrepancies were resolved through discussion until consensus was reached.
V-B Limitations
One limitation of WIRL is that it recommends substitutions for unresolved elements using only predefined variables available in the surrounding context. However, in some cases, the necessary variables may not exist in the current context, and declaring new variables becomes necessary. Since WIRL is specifically designed to support context-aware code adaptation, handling variable declarations falls outside its current scope. We plan to extend WIRL to address such scenarios in future work.
Another limitation is that WIRL currently supports only the Java programming language and has been evaluated solely on Java-based code. This is partly due to the limited availability of cross-language datasets involving real-world code reuse, such as from Stack Overflow to GitHub. In future research, it would be valuable to explore how to generalize and adapt WIRL to support copy-paste-modify practices in other programming languages.
VI Related Work
VI-A Code Snippet Adaptation
Numerous empirical studies and tools have been proposed to tackle code snippet adaptation. Wu et al. [48] conducted the first empirical investigation into how developers reuse Stack Overflow code, identifying five reuse patterns: exact copy, cosmetic modification, non-cosmetic modification (e.g., adding/removing statements or altering structure), idea conversion (writing from scratch), and information use. Extending this work, Chen et al. [47] examined how Stack Overflow answers are reused in GitHub projects. Using a hybrid of clone detection, keyword search, and manual inspection, they identified genuine reuse cases and proposed 11 sub-types of non-cosmetic modifications. They found that most developers reused at least some code, frequently applying statement-level or structural edits.
Zhang et al.[46] explored context-based adaptations through interviews with 21 developers, highlighting their prevalence and significance. They proposed four adaptation patterns and later introduced an interactive prompt engineering approach[30], involving executor and evaluator roles. Though iterative, this approach is not agent-based, as it cannot dynamically interact with surrounding context. Key differences from WIRL include method-level (vs. block-level) adaptation and distinct strategies for context acquisition.
Initial efforts to automate variable resolution include CSNIPPEX[29] and NLP2Testable[27], which declare default-valued variables to resolve missing references. While these address immediate issues, they lack the capacity for dynamic context adaptation, making them unsuitable for complex reuse scenarios. As Zhang et al. [46] noted, 56.1% of reuse cases require context-based adaptation, underscoring the need for more robust solutions like WIRL.
Some approaches leverage code clones and templates. Wightman et al.’s SniptMatch [24], an Eclipse plugin, maps natural language queries to snippets with variable placeholders, which users must manually fill. However, its focus is on preventing syntax errors, not enabling active adaptation. Zhang et al. [28] proposed ExampleStack, a Chrome extension that integrates online examples by leveraging past adaptations from open-source projects. It highlights frequently modified regions while preserving unchanged parts, but its dependency on existing adaptations limits its generalization to unseen cases. Lin et al. introduced CCDemon[25], which recommends edits to pasted code based on local clones, contextual elements, and naming conventions. They later proposed MICoDe[22], an Eclipse plugin for summarizing reusable module-level patterns. Similarly, EUKLAS [19] for JavaScript automatically imports missing variables and functions from the local project. While effective, these tools are constrained to local contexts and struggle with adaptation across different projects or unseen scenarios.
Another line of work treats adaptation as a deobfuscation task [20, 21], masking variables and recovering them via context-aware models pre-trained with dataflow anonymization objectives. Though conceptually aligned, these methods are computationally inefficient for scenarios involving only a few unresolved variables and are typically limited to single-file contexts, making them inadequate for broader project-level adaptation.
Other approaches support adaptation in alternative scenarios. JigSaw[26] evaluates the compatibility of inserted code via AST analysis to assist in integrating missing functionality. APIzation[23] helps convert code snippets into well-defined method declarations by statically analyzing parameters and return values. While related, these methods target different application goals than WIRL.
In contrast, WIRL enables automatic snippet adaptation with a particular emphasis on resolving unresolved variables through dynamic context analysis and LLM-powered recommendations.
VI-B LLM-based Agents for Software Engineering
LLM-based agents have become transformative tools in automating software engineering tasks through iterative reasoning and code synthesis [34, 37]. Recent work highlights their versatility: Huang et al.[31] introduced AgentCoder, a multi-agent system for code generation via dynamic testing; Lin et al.[33] modeled software processes (e.g., Scrum) by assigning LLMs to development roles (e.g., architect, tester); Zhang et al.[39] and Yang et al.[38] developed agents for test generation using code navigation toolchains. In code maintenance, Batole et al.’s LocalizeAgent[35] recommends refactorings, while Bouzenia et al.’s RepairAgent[40] performs program repair via dynamic prompt engineering. Tao et al.[36] automated GitHub issue resolution with role-specialized agents, and Ke et al.[32] applied LLM agents to flaky test repair.
Despite these advancements, no prior work addresses context-aware code wiring, specifically, automatically resolving unresolved variables by substituting them with existing ones. WIRL fills this gap as the first framework to integrate agent-based planning, customized toolkit, and dynamically updated prompts for automated identifier harmonization.
VII Conclusions and Future Work
In this paper, we introduced WIRL, a novel LLM-based agent designed to facilitate context-aware code wiring, an essential but underexplored aspect of the copy-paste-modify paradigm in software development. Unlike prior approaches that rely on static heuristics or historical templates, WIRL keeps a dynamically updated prompt to collect and record useful context information through a hybrid architecture that combines structured tool invocation with autonomous reasoning capabilities of LLMs. By formulating the adaptation task as an RAG infilling problem, WIRL aligns closely with the natural strengths of LLMs in code completion. Our evaluation on a carefully curated benchmark dataset demonstrates that WIRL achieves state-of-the-art performance, significantly outperforming both commercial IDEs and advanced LLMs in exact match precision and recall. These results underscore the practical value of intelligent, context-sensitive tooling for reducing developer effort and improving the reliability of reused code.
Looking ahead, we envision WIRL as a stepping stone toward more general-purpose adaptation agents that can support a broader range of integration and transformation tasks. We believe this direction holds strong potential for reshaping how developers interact with reused and AI-generated code in modern software engineering workflows.
VIII Data Availability
The replication package, including tools and data, is publicly available in [8].